OpenStudy (anonymous):

Can someone tell me if I'm correct? The property of polynomial multiplication that allows us to combine like terms like this: 2x^3 + 5x^3 = 7x^3 is ASSOSCIATIVE.

OpenStudy (tkhunny):

Hmmm.... No. You have changed no association. 2x^3 + 5x^3 = (2+5)x^3 Look familiar, yet?

OpenStudy (anonymous):

Commutative?

OpenStudy (tkhunny):

This is commutative: a*b*c = a*c*b -- See how we changed the order of things? This is associative: (a*b)*c = a*(b*c) -- See how we just grouped them differently? See also how there is ONLY multiplication going on, here? This is commutative: a + b + c = a + c + b -- See how we changed the order of things? This is associative: (a+b)+c = a+(b+c) -- See how we just grouped them differently? See also how there is ONLY addition going on, here? You have a property that involves both multiplication AND addition.

OpenStudy (anonymous):

I get it. So then it has to be distributive?

OpenStudy (tkhunny):

Yes! The Distributive Property of Multiplication over Addition. Sometimes it helps to use the whole name, albeit quite tedious.

OpenStudy (anonymous):

Alright. I understand it now. Thank you very much